# Diffusing States and Matching Scores: A New Framework for Imitation Learning

### v1

@@ Abstract
Adversarial Imitation Learning is traditionally framed as a two-player zero-sum game between a learner and an adversarially chosen cost function, and can therefore be thought of as the sequential generalization of a Generative Adversarial Network (GAN). However, in recent years, diffusion models have emerged as a non-adversarial alternative to GANs that merely require training a score function via regression, yet produce generations of a higher quality. In response, we investigate how to lift insights from diffusion modeling to the sequential setting. We propose diffusing states and performing score-matching along diffused states to measure the discrepancy between the expert’s and learner’s states. Thus, our approach only requires training score functions to predict noises via standard regression, making it significantly easier and more stable to train than adversarial methods. Theoretically, we prove first- and second-order instance-dependent bounds with linear scaling in the horizon, proving that our approach avoids the compounding errors that stymie offline approaches to imitation learning. Empirically, we show our approach outperforms both GAN-style imitation learning baselines and discriminator-free imitation learning baselines across various continuous control problems, including complex tasks like controlling humanoids to walk, sit, crawl, and navigate through obstacles.
@@ 1 Introduction
Fundamentally, in imitation learning (IL, Osa et al. (2018)), we want to match the sequential behavior of an expert demonstrator. Different notions of what matching should mean for IL have been proposed in the literature, from $f$-divergences (Ho & Ermon, 2016; Ke et al., 2021) to Integral Probability Metrics (IPMs, Müller (1997); Sun et al. (2019); Kidambi et al. (2021); Swamy et al. (2021); Chang et al. (2021); Song et al. (2024)). To compute the chosen notion of divergence from the expert demonstrations so that the learner can then optimize it, it is common to train a discriminator (i.e. a classifier) between expert and learner data. This discriminator is then used as a reward function for a policy update, an approach known as inverse reinforcement learning (IRL, Abbeel & Ng (2004); Ziebart et al. (2008)). Various losses for training discriminators have been proposed (Ho & Ermon, 2016; Kostrikov et al., 2018; Fu et al., 2017; Ke et al., 2021; Swamy et al., 2021; Chang et al., 2024), and IRL has been applied in real life in domains like routing (Barnes et al., 2023) and LLM fine-tuning (Wulfmeier et al., 2024).
As observed by Finn et al. (2016), inverse RL can be seen as the sequential generalization of a Generative Adversarial Network (GAN) (Goodfellow et al., 2020). However, even in the single-shot setting, GANs are known to suffer from issues like unstable training dynamics and mode collapse  (Miyato et al., 2018; Arjovsky & Bottou, 2017). In contrast, Score-Based Diffusion Models (Ho et al., 2020; Song & Ermon, 2019; Song et al., 2020) are known to be more stable to train and produce higher quality samples in domains like audio and video (Rombach et al., 2022; Ramesh et al., 2022; Kong et al., 2020). Intuitively, diffusion models corrupt samples from a target distribution with noise and train a generative model to reverse this corruption. Critically, these models are trained via score-matching: a non-adversarial, purely regression-based procedure in which a network is trained to match the score (i.e., gradient of the log probability) of the target distribution.
@@
We demonstrate the advantages of our framework in both theory and practice. In theory, we show that our approach can achieve first- and second-order instance-dependent regret bounds, as well as a linear scaling in the horizon to model-misspecification errors arising from expert not being realizable by the learner’s policy class or potential optimization error in score-matching and policy search. Thus, we establish that
By lifting score-matching to IRL, SMILING provably avoids compounding errors and achieves instance-dependent bounds.
Intuitively, the second-order bounds mean that the performance gap between our learned policy and the expert automatically shrinks when either the expert or the learned policy has low variance in terms of their performance under the ground-truth reward (e.g. when the expert and dynamics are relatively deterministic). This is because when we perform score-matching, we are actually minimizing the squared Hellinger distance between the learner and the expert’s state distributions, which has shown to play an important role in achieving second-order regret bounds in the Reinforcement Learning setting (Wang et al., 2024c). The ability to achieve instance-dependent bounds demonstrates the theoretical benefit of Diffusion Score divergence over other metrics such as IPMs and $f$-divergences. In practice, we show that under in the IL from observation only setting (Torabi et al., 2018; Sun et al., 2019), SMILING outperforms adversarial GAN-based IL baselines, discriminator-free IL baselines, and Behavioral Cloning [^1] This is despite the fact that BC requires the expert actions, while we do not. on complex tasks such as controlling humanoids to walk, sit, crawl, and navigate through poles (see Figure 1). This makes SMILING the first IRL method to solve multiple tasks on the recently released HumanoidBench benchmark (Sferrazza et al., 2024) using only the state information of the expert demonstrations. We release the code for all experiments at https://github.com/ziqian2000/SMILING.
@@ 2 Related Works
Inverse Reinforcement Learning. Starting with the seminal work of Abbeel & Ng (2004), various authors have proposed solving the problem of imitation via inverse RL (Ziebart et al., 2008; Ratliff et al., 2006; Sun et al., 2019; Kidambi et al., 2021; Chang et al., 2021). As argued by Swamy et al. (2021), these approaches can be thought of as minimizing an integral probability metric (IPM) between learner and expert behavior via an adversarial training procedure – we refer interested readers to their paper for a full list. Many other IRL approaches can instead be framed as minimizing an $f$-divergence via adversarial training (Ke et al., 2021), including GAIL (Ho & Ermon, 2016), DAC (Kostrikov et al., 2018), AIRL (Fu et al., 2017), FAIRL (Ghasemipour et al., 2020), and f-GAIL (Zhang et al., 2020). However, all of the preceding approaches involve training a discriminator, while our technique only requires fitting score functions and does not seek to approximate either an $f$-divergence or an IPM. Wang et al. (2024a) also explored the use of diffusion models for imitation learning. They insert the score matching loss into the $f$-divergence objective of the discriminator. Thus, their method also belongs to the $f$-divergence framework. Huang et al. (2024) also utilize diffusion models for imitation learning, proposing to train a conditional diffusion model as a discriminator for single-step state transitions between the expert and the current policy. Their approach remains within the classic $f$-divergence framework as well. Additionally, various authors have proposed using techniques to either stabilize this training procedure (e.g. via boosting, Chang et al. (2024)) or reducing the amount of interaction required during the RL step (Swamy et al., 2023; Ren et al., 2024; Sapora et al., 2024) – as we focus on improving the reward function used in inverse RL, our approach is orthogonal to and could be naturally applied on top of these techniques.
@@
Discriminator-Free Inverse Reinforcement Learning. SQIL (Reddy et al., 2019) replaces training a discriminator with a fixed reward function ($+1$ for expert data, 00 for learner data). Unfortunately, this means that a performant learner is dis-incentivized to perform expert-like behavior, leading to dramatic drops in performance (Barde et al., 2020). The AdRIL algorithm of Swamy et al. (2021) uses techniques from functional gradients to address this issue, but requires the use off an off-policy RL algorithm to implement, while our framework makes no such assumptions. ASAF (Barde et al., 2020) proposes using the prior policy to compute the optimal $f$-divergence discriminator in closed form, while we focus on a different class of divergences. IQ-Learn (Garg et al., 2021) proposes to perform IRL in the space $Q$ functions, but can therefore suffer from poor performance on problems with stochastic dynamics (Ren et al., 2024) or when the $Q$-function is more complex than the reward function (e.g. navigating to a goal in a maze). Jain et al. (2024) propose measuring differences between learner and expert behavior in terms of successor features, but are only able to optimize over deterministic policies. DRIL (Brantley et al., 2019) relies on ensemble methods and uses the disagreement among models in the ensemble as a surrogate cost function, making it challenging to prove similar theoretical guarantees to our technique.
In contrast, our method instead trains score functions on the state distribution of the policy via the standard score matching objective. We show that our approach can require less expressive function approximators than discriminator-based methods (see Section 5.1). Additionally, our approach achieves solid first- and second-order sample complexity bound and also delivers strong empirical performance on challenging humanoid control tasks.
@@
Diffusion Models for Decision Making. Diffusion models have been widely studied in the context robot learning, where diffusion models have demonstrated a strong ability to capture complex and multi-modal data. Recent works have explored using diffusion models to perform behavioral cloning   Chi et al. (2023); Pearce et al. (2023); Block et al. (2023); Team et al. (2024) as well as using diffusion models to capture trajectory-level dynamics  (Janner et al., 2022; Ajay et al., 2022). These prior work exclusively rely on offline data for training. We instead focus on lifting insights from diffusion modeling and score matching to IRL, which uses both offline expert data and also online interaction with the environment. Unlike prior diffusion BC work, we use diffusion models for both the expert’s and the learner’s state distributions, not just the expert’s action-conditional distributions. This means we do not require expert action labels, which can be challenging to acquire in practice.
@@ 3.2 Imitation Learning
In imitation learning, we are usually given expert demonstration in the form of state-action-next-state tuples, and the goal is to learn a policy that mimics the expert behavior. We do not have access to the ground-truth cost function $c^{\star}$. In this paper, we consider a harder setting where only states are given. Specifically, given a dataset $\mathcal{D}^{e}=\{s^{(i)}\}_{i=1}^{N}$ of expert demonstrations, the goal is to learn a policy $\pi$ that minimizes the discrepancy between the state distributions of policy $\pi$ and the expert policy $\pi^{e}$. In the inverse reinforcement learning setting, in addition to the expert demonstrations $\mathcal{D}^{e}$, we can interact with the environment to collect more (reward-free) learner trajectories.
@@ 3.3 Diffusion Models
Generally speaking, the forward process of a diffusion model adds noise to a sample from the data distribution $p_{0}$. We can formalize this via a stochastic differential equation (SDE). For simplicity in this paper, we consider the Ornstein-Uhlenbeck (OU) process:
$\mathop{}\mathrm{d}\mathbf{x}_{t}=-\mathbf{x}_{t}\mathop{}\mathrm{d}t+\sqrt{2}\mathop{}\mathrm{d}B_{t},\quad\mathbf{x}_{0}\sim p_{0},$ (1)
where $B_{t}$ is the standard Brownian motion in $\mathbb{R}^{d}$. It is known that the reverse process $(\mathbf{y}_{t})_{t\in[0,T]}$ satisfies the following reverse-time SDE (Anderson, 1982):
$\mathop{}\mathrm{d}\mathbf{y}_{t}=(\mathbf{y}_{t}+2\nabla\log p_{T-t}(\mathbf{y}_{t}))\mathop{}\mathrm{d}t+\sqrt{2}\mathop{}\mathrm{d}B_{t},\quad\mathbf{y}_{0}\sim p_{T}$
where $U(T)$ denotes the uniform distribution on $[0,T]$, and $B_{t}$ now denotes reversed Brownian motion. The gradient of the log probability $\nabla\log p_{t}$ is called the score function of the distribution $p_{t}$. When the score function is known, we can use it to generate samples from $p_{0}$ by simulating the reverse process. Estimating the score function $\nabla\log p_{t}$ is called score matching, which is typically done via minimizing the following regression-based loss:
$\min_{g}\mathop{\mathbb{E}}_{\mathbf{x}\sim p_{0}}\mathop{\mathbb{E}}_{t\sim U(T)}\mathop{\mathbb{E}}_{\mathbf{x}_{t}\sim q_{t}(\cdot\,|\,\mathbf{x})}\Big{\|}g(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}\,|\,\mathbf{x})\Big{\|}_{2}^{2}.$
We use $q_{t}(\mathbf{x}_{t}\,|\,\mathbf{x})$ to denote the conditional distribution at time $t$ of the forward process conditioned on the initial state $\mathbf{x}$, which has closed form $q_{t}(\mathbf{x}_{t}\,|\,\mathbf{x})=\mathcal{N}(\mathbf{x}e^{-t},(1-e^{-2t})I)$.
Applying Diffusion to State Distributions. For a policy $\pi$, we will use $p_{t}^{\pi}$ to denote the marginal distribution at time $t$ obtained by applying the forward diffusion process to $d^{\pi}$, i.e., initial samples are drawn from $p_{0}^{\pi}\coloneqq d^{\pi}$. When presenting asymptotic results, we will use $O(\cdot)$ to hide constants and $\widetilde{O}(\cdot)$ to hide constants and logarithmic factors.
@@ 4 Algorithm
4:      Update policy $\pi^{(k)}$ via RL (e.g., SAC) on cost $c^{(k)}$ (Eq. 4)
$\pi^{(k)}\leftarrow\text{RL}(c^{(k)})$
@@
We begin by introducing a novel discrepancy measure between two distributions that will be leveraged by algorithm:
@@
Figure 2(a) illustrates the DS divergence. It measures the difference between two distributions by comparing their diffusion score functions within one diffusion process. It is analogous to Fisher divergence but differs by incorporating an expectation over the diffusion process. DS divergence is a strong divergence in the sense that, whenever the DS divergence between the two distributions is small, the KL divergence, Hellinger distance, and total variation distance are all small (see Lemma 2 and also Chen et al. (2022); Oko et al. (2023)).
@@
(a) Illustration of DS Divergence (Definition 1).
(b) Illustration of SMILING (Algorithm 1).
Figure 2: Figure (a): The two curves represent the forward diffusion process of distributions $P$ and $Q$. DS Divergence measures the squared difference between the diffusion score functions, $\nabla\log P_{t}(s_{t})$ and $\nabla\log Q_{t}(s_{t})$, along the forward diffusion process of $P$. Figure (b): SMILING first pre-trains a diffusion model from the expert’s data. It then iteratively trains diffusion models on learner’s data and performs RL to optimize a cost function formed by the learner’s score function and the pre-trained expert score function. The cost function is designed to faithfully approximate the DS divergence between the learner and the expert.
@@
where we have approximated the expert’s diffusion score function by $g^{e}(s_{t},t)$. However, $\ell(\pi)$ is not directly computable since we do not know the learner’s score function $\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$. A naive approach would be to directly learn an estimator for $\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$ via score matching and substitute it into the equation. However, as we prove in Appendix A, this can introduce significant errors due to the unboundedness of a difference of score functions and variance-related concerns. We now derive a method that does not suffer from this issue.
Recall that we defined $q_{t}(s_{t}\,|\,s)$ as the distribution of $s_{t}$ given the initial sample $s$, which is a simple Gaussian distribution with an appropriately scaled variance. To develop our method, we first note that, luckily, $\nabla_{s_{t}}\log q_{t}(s_{t}|s)$ is an unbiased estimator of the score $\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$, because $\mathop{\mathbb{E}}_{s|s_{t}}[\nabla_{s_{t}}\log q_{t}(s_{t}|s)]=\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$ (Song et al., 2020). Given this, perhaps the most immediate strategy would be to simply replace $\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$ by its unbiased estimator, $\nabla_{s_{t}}\log q_{t}(s_{t}|s)$, in Eq. 2. However, in contrast to linear objectives like an IPM, to approximate a squared objective like the DS Divergence accurately, we need to get both the expectation (i.e. have an unbiased estimator) as well as the variance correct. If we don’t, then $\mathop{\mathbb{E}}_{s\sim d^{\pi},t\sim U(T),s_{t}\sim q_{t}(\cdot|s)}\|g^{e}(s_{t},t)-\nabla_{s_{t}}\log q_{t}(s_{t}|s)\|_{2}^{2}$ will differ from $\ell(\pi)$ by a term related to variance of the estimator $\nabla_{s_{t}}\log q_{t}(s_{t}|s)$, i.e.
@@
$\mathcal{L}(\pi,g)$ $:=\mathop{\mathbb{E}}_{s\sim d^{\pi},t\sim U(T),s_{t}\sim q_{t}(\cdot|s)}\left\|g^{e}(s_{t},t)-\nabla_{s_{t}}\log q_{t}(s_{t}|s)\right\|_{2}^{2}$
$\qquad-\mathop{\mathbb{E}}_{s\sim d^{\pi},t\sim U(T),s_{t}\sim q_{t}(\cdot|s)}\left\|g(s_{t},t)-\nabla_{s_{t}}\log q_{t}(s_{t}|s)\right\|_{2}^{2}.$
Then, minimizing ^$\hat{\ell}(\pi)$ is equivalent to solving the two-player zero sum game $\min_{\pi}\max_{g}\mathcal{L}(\pi,g).$ To solve this game, we propose following a no-regret strategy over $g$ against a best response over $\pi$ (i.e. a dual algorithm (Swamy et al., 2021)). Specifically, Algorithm 1 applies Follow-the-Leader (Shalev-Shwartz et al., 2012)) to optimize $g$ in Line 3, [^4] Note that $\mathcal{L}(\pi,g)$ is a square-loss functional with respect to $g$. Since square loss is strongly convex, FTL is no-regret. In contrast, when optimizing an IPM, one has to use Follow-the-Regularized-Leader (FTRL) to have the no-regret property (e.g., Sun et al. (2019); Swamy et al. (2021)), making implementation more complicated. and performs best response computation over $\pi$ via RL (Line 4) under cost function $c^{k}(s)$:

### v2

@@ Abstract
Adversarial Imitation Learning is traditionally framed as a two-player zero-sum game between a learner and an adversarially chosen cost function, and can therefore be thought of as the sequential generalization of a Generative Adversarial Network (GAN). However, in recent years, diffusion models have emerged as a non-adversarial alternative to GANs that merely require training a score function via regression, yet produce generations of higher quality. In response, we investigate how to lift insights from diffusion modeling to the sequential setting. We propose diffusing states and performing score-matching along diffused states to measure the discrepancy between the expert’s and learner’s states. Thus, our approach only requires training score functions to predict noises via standard regression, making it significantly easier and more stable to train than adversarial methods. Theoretically, we prove first- and second-order instance-dependent bounds with linear scaling in the horizon, proving that our approach avoids the compounding errors that stymie offline approaches to imitation learning. Empirically, we show our approach outperforms both GAN-style imitation learning baselines and discriminator-free imitation learning baselines across various continuous control problems, including complex tasks like controlling humanoids to walk, sit, crawl, and navigate through obstacles.
@@ 1 Introduction
Fundamentally, in imitation learning (IL, Osa et al. (2018)), we want to match the sequential behavior of an expert demonstrator. Different notions of what matching should mean for IL have been proposed in the literature, from $f$-divergences (Ho & Ermon, 2016; Ke et al., 2021) to Integral Probability Metrics (IPMs, Müller (1997); Sun et al. (2019); Kidambi et al. (2021); Swamy et al. (2021); Chang et al. (2021); Song et al. (2024)). To compute the chosen notion of divergence from the expert demonstrations so that the learner can then optimize it, it is common to train a discriminator (i.e. a classifier) between expert and learner data. This discriminator is then used as a reward function for a policy update, an approach known as inverse reinforcement learning (IRL, Abbeel & Ng (2004); Ziebart et al. (2008)). Various losses for training discriminators have been proposed (Ho & Ermon, 2016; Kostrikov et al., 2019; Fu et al., 2017; Ke et al., 2021; Swamy et al., 2021; Chang et al., 2024), and IRL has been applied in real life in domains like routing (Barnes et al., 2023) and LLM fine-tuning (Wulfmeier et al., 2024).
As observed by Finn et al. (2016), inverse RL can be seen as the sequential generalization of a Generative Adversarial Network (GAN) (Goodfellow et al., 2020). However, even in the single-shot setting, GANs are known to suffer from issues like unstable training dynamics and mode collapse  (Miyato et al., 2018; Arjovsky & Bottou, 2017). In contrast, Score-Based Diffusion Models (Ho et al., 2020; Song & Ermon, 2019; Song et al., 2021) are known to be more stable to train and produce higher quality samples in domains like audio and video (Rombach et al., 2022; Ramesh et al., 2022; Kong et al., 2021). Intuitively, diffusion models corrupt samples from a target distribution with noise and train a generative model to reverse this corruption. Critically, these models are trained via score-matching: a non-adversarial, purely regression-based procedure in which a network is trained to match the score (i.e., gradient of the log probability) of the target distribution.
@@
We demonstrate the advantages of our framework in both theory and practice. In theory, we show that our approach can achieve first- and second-order instance-dependent regret bounds, as well as a linear scaling in the horizon to model-misspecification errors arising from expert not being realizable by the learner’s policy class or potential optimization error in score-matching and policy search. Thus, we establish that, by lifting score-matching to IRL, SMILING provably avoids compounding errors and achieves instance-dependent bounds. Intuitively, the second-order bounds mean that the performance gap between our learned policy and the expert automatically shrinks when either the expert or the learned policy has low variance in terms of their performance under the ground-truth reward (e.g. when the expert and dynamics are relatively deterministic). This is because when we perform score-matching, we are actually minimizing the squared Hellinger distance between the learner and the expert’s state distributions, which has shown to play an important role in achieving second-order regret bounds in the Reinforcement Learning setting (Wang et al., 2024c). The ability to achieve instance-dependent bounds demonstrates the theoretical benefit of Diffusion Score divergence over other metrics such as IPMs and $f$-divergences. In practice, we show that under in the IL from observation only setting (Torabi et al., 2018; Sun et al., 2019), SMILING outperforms adversarial GAN-based IL baselines, discriminator-free IL baselines, and Behavioral Cloning [^1] This is despite the fact that BC requires expert actions, while we do not. on complex tasks such as controlling humanoids to walk, sit, crawl, and navigate through poles (see Figure 3). This makes SMILING the first IRL method to solve multiple tasks on the recently released HumanoidBench benchmark (Sferrazza et al., 2024) using only the state information of the expert demonstrations. We release the code for all experiments at https://github.com/ziqian2000/SMILING.
@@ 2 Related Works
Inverse Reinforcement Learning. Starting with the seminal work of Abbeel & Ng (2004), various authors have proposed solving the problem of imitation via inverse RL (Ziebart et al., 2008; Ratliff et al., 2006; Sun et al., 2019; Kidambi et al., 2021; Chang et al., 2021). As argued by Swamy et al. (2021), these approaches can be thought of as minimizing an integral probability metric (IPM) between learner and expert behavior via an adversarial training procedure – we refer interested readers to their paper for a full list. Many other IRL approaches can instead be framed as minimizing an $f$-divergence via adversarial training (Ke et al., 2021), including GAIL (Ho & Ermon, 2016), DAC (Kostrikov et al., 2019), AIRL (Fu et al., 2017), FAIRL (Ghasemipour et al., 2020), f-GAIL (Zhang et al., 2020), and f-IRL (Ni et al., 2021). However, all of the preceding approaches involve training a discriminator, while our technique only requires fitting score functions and does not seek to approximate either an $f$-divergence or an IPM. Lai et al. (2024); Wang et al. (2024a) also explored the use of diffusion models for imitation learning. They insert the score matching loss into the $f$-divergence objective of the discriminator. Thus, their method also belongs to the $f$-divergence framework. Huang et al. (2024) also utilize diffusion models for imitation learning, proposing to train a conditional diffusion model as a discriminator for single-step state transitions between the expert and the current policy. Their approach remains within the classic $f$-divergence framework as well. Additionally, various authors have proposed using techniques to either stabilize this training procedure (e.g. via boosting, Chang et al. (2024)) or reducing the amount of interaction required during the RL step (Swamy et al., 2023; Ren et al., 2024; Sapora et al., 2024) – as we focus on improving the reward function used in inverse RL, our approach is orthogonal to and could be naturally applied on top of these techniques.
@@
Discriminator-Free Inverse Reinforcement Learning. SQIL (Reddy et al., 2020) replaces training a discriminator with a fixed reward function ($+1$ for expert data, 00 for learner data). Unfortunately, this means that a performant learner is dis-incentivized to perform expert-like behavior, leading to dramatic drops in performance (Barde et al., 2020). The AdRIL algorithm of Swamy et al. (2021) uses techniques from functional gradients to address this issue, but requires the use off an off-policy RL algorithm to implement, while our framework makes no such assumptions. ASAF (Barde et al., 2020) proposes using the prior policy to compute the optimal $f$-divergence discriminator in closed form, while we focus on a different class of divergences. IQ-Learn (Garg et al., 2021) proposes to perform IRL in the space $Q$ functions, but can therefore suffer from poor performance on problems with stochastic dynamics (Ren et al., 2024) or when the $Q$-function is more complex than the reward function (e.g. navigating to a goal in a maze). Al-Hafez et al. (2023) propose to use a reward regularization to overcome the instabilities of prior work that learns $Q$-functions. Sikchi et al. (2024a) point out that prior methods rely on restrictive coverage assumptions and propose a new method that learns to imitate from arbitrary off-policy data. Jain et al. (2024) propose measuring differences between learner and expert behavior in terms of successor features, but are only able to optimize over deterministic policies. Sikchi et al. (2024b) suggest to learn a multi-step utility function that quantifies the impact of actions on the agent’s divergence from the expert’s visitation distribution. DRIL (Brantley et al., 2019) relies on ensemble methods and uses the disagreement among models in the ensemble as a surrogate cost function, making it challenging to prove similar theoretical guarantees to our technique. Dadashi et al. (2021) frame IL as minimizing the Wasserstein distance in its primal form rather than the dual commonly used in prior discriminator-based works. Liu et al. (2021) also consider score matching but they use energy-based models instead of diffusion models.
@@
Diffusion Models for Decision Making. Diffusion models have been widely studied in the context of robot learning, where diffusion models have demonstrated a strong ability to capture complex and multi-modal data. Recent works have explored using diffusion models to perform behavioral cloning (Chi et al., 2023; Pearce et al., 2023; Chen et al., 2024; Block et al., 2023; Team et al., 2024) and for reward guidance (Nuti et al., 2024) as well as using diffusion models to capture trajectory-level dynamics  (Janner et al., 2022; Ajay et al., 2022). We instead focus on lifting insights from diffusion modeling and score matching to IRL, which uses both offline expert data and also online interaction with the environment. Unlike prior diffusion BC work, we use diffusion models for both the expert’s and the learner’s state distributions, not just the expert’s action-conditional distributions. This means we do not require expert action labels, which can be challenging to acquire in practice.
@@ 3.2 Imitation Learning
In imitation learning, we are usually given expert demonstration in state-action-next-state tuples and aim to learn a policy $\pi$ that mimics the expert policy $\pi^{e}$ without access to the ground-truth cost function $c^{\star}$. This paper considers a harder setting where only states are given. Specifically, given a dataset of state demonstrations $\mathcal{D}^{e}=\{s^{(i)}\}_{i=1}^{N}$ sampled from expert, we aim to learn a policy $\pi$ that minimizes the discrepancy between its state distributions and the expert’s. In addition, we can interact with the environment to collect more (reward-free) learner trajectories.
@@ 3.3 Diffusion Models
The forward process of a diffusion model adds noise to a sample from the data distribution $p_{0}$. We can formalize it via a stochastic differential equation (SDE). For simplicity in this paper, we consider the Ornstein-Uhlenbeck (OU) process: $\mathop{}\mathrm{d}\mathbf{x}_{t}=-\mathbf{x}_{t}\mathop{}\mathrm{d}t+\sqrt{2}\mathop{}\mathrm{d}B_{t}$, $\mathbf{x}_{0}\sim p_{0}$, where $B_{t}$ is the standard Brownian motion in $\mathbb{R}^{d}$. It is known that the reverse process $(\mathbf{y}_{t})_{t\in[0,T]}$ satisfies the following reverse-time SDE (Anderson, 1982): $\mathop{}\mathrm{d}\mathbf{y}_{t}=(\mathbf{y}_{t}+2\nabla\log p_{T-t}(\mathbf{y}_{t}))\mathop{}\mathrm{d}t+\sqrt{2}\mathop{}\mathrm{d}B_{t},\mathbf{y}_{0}\sim p_{T}$ where $U(T)$ denotes the uniform distribution on $[0,T]$, and $B_{t}$ now denotes reversed Brownian motion. The gradient of the log probability $\nabla\log p_{t}$ is called the score function of the distribution $p_{t}$. When the score function is known, we can use it to generate samples from $p_{0}$ by simulating the reverse process. Estimating the score function $\nabla\log p_{t}$ is called score matching, which is typically done via minimizing the regression-based loss: $\min_{g}\mathop{\mathbb{E}}_{\mathbf{x}\sim p_{0}}\mathop{\mathbb{E}}_{t\sim U(T)}\mathop{\mathbb{E}}_{\mathbf{x}_{t}\sim q_{t}(\cdot\,|\,\mathbf{x})}\|g(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}\,|\,\mathbf{x})\|_{2}^{2}.$ We use $q_{t}(\mathbf{x}_{t}\,|\,\mathbf{x})$ to denote the conditional distribution at time $t$ of the forward process conditioned on the initial state $\mathbf{x}$, which has closed form $q_{t}(\mathbf{x}_{t}\,|\,\mathbf{x})=\mathcal{N}(\mathbf{x}e^{-t},(1-e^{-2t})I)$.
Applying Diffusion to State Distributions. For a policy $\pi$, we use $p_{t}^{\pi}$ to denote the marginal distribution at time $t$ obtained by applying the forward diffusion process to $d^{\pi}$, i.e., initial samples are drawn from $p_{0}^{\pi}\coloneqq d^{\pi}$. When presenting asymptotic results, we use $O(\cdot)$ to hide constants and $\widetilde{O}(\cdot)$ to hide constants and logarithmic factors.
@@ 4 Algorithm
(a) Illustration of DS Divergence (Definition 1).
(b) Illustration of SMILING (Algorithm 1).
Figure 1: Figure (a): The two curves represent the forward diffusion process of distributions $P$ and $Q$. DS Divergence measures the squared difference between the diffusion score functions, $\nabla\log P_{t}(s_{t})$ and $\nabla\log Q_{t}(s_{t})$, along the forward diffusion process of $P$. Figure (b): SMILING first pre-trains a diffusion model from the expert’s data. It then iteratively trains diffusion models on learner’s data and performs RL to optimize a cost function formed by the learner’s score function and the pre-trained expert score function. The cost function is designed to faithfully approximate the DS divergence (Definition 1) between the learner and the expert.
@@
4:      Update policy $\pi^{(k)}$ via RL (e.g., SAC) on cost $c^{(k)}$ (Eq. 3): $\pi^{(k)}\leftarrow\text{RL}(c^{(k)})$
@@
We introduce a novel discrepancy measure between two distributions that our algorithm leverages:
@@
Figure 1(a) illustrates the DS divergence. It measures the difference between two distributions by comparing their diffusion score functions within one diffusion process. It is analogous to Fisher divergence but differs by incorporating an expectation over the diffusion process. DS divergence is a strong divergence in the sense that, whenever the DS divergence between the two distributions is small, the KL divergence, Hellinger distance, and total variation distance are all small (see Lemma 2 and also Chen et al. (2023b); Oko et al. (2023)).
@@
where we have approximated the expert’s diffusion score function by $g^{e}(s_{t},t)$. However, $\ell(\pi)$ is not directly computable since we do not know the learner’s score function $\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$. A naive approach would be to directly learn an estimator for $\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$ via score matching and substitute it into the equation. However, as we prove in Appendix B, this can introduce significant errors due to the unboundedness of a difference of score functions and variance-related concerns. We now derive a method that does not suffer from this issue.
Recall that we defined $q_{t}(s_{t}\,|\,s)$ as the distribution of $s_{t}$ given the initial sample $s$, which is a simple Gaussian distribution with an appropriately scaled variance. To develop our method, we first note that, luckily, $\nabla_{s_{t}}\log q_{t}(s_{t}|s)$ is an unbiased estimator of the score $\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$, because $\mathop{\mathbb{E}}_{s|s_{t}}[\nabla_{s_{t}}\log q_{t}(s_{t}|s)]=\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$ (Song et al., 2021). Given this, perhaps the most immediate strategy would be to simply replace $\nabla_{s_{t}}\log p^{\pi}_{t}(s_{t})$ by its unbiased estimator, $\nabla_{s_{t}}\log q_{t}(s_{t}|s)$, in Eq. 1. However, in contrast to linear objectives like an IPM, to approximate a squared objective like the DS Divergence accurately, we need to get both the expectation (i.e. have an unbiased estimator) as well as the variance correct. If we don’t, then $\mathop{\mathbb{E}}_{s\sim d^{\pi},t\sim U(T),s_{t}\sim q_{t}(\cdot|s)}\|g^{e}(s_{t},t)-\nabla_{s_{t}}\log q_{t}(s_{t}|s)\|_{2}^{2}$ will differ from $\ell(\pi)$ by a term related to variance of the estimator $\nabla_{s_{t}}\log q_{t}(s_{t}|s)$, i.e.
@@
$\mathcal{L}(\pi,g)$ $:=\mathop{\mathbb{E}}_{s\sim d^{\pi},t\sim U(T),s_{t}\sim q_{t}(\cdot|s)}\left(\left\|g^{e}(s_{t},t)-\nabla_{s_{t}}\log q_{t}(s_{t}|s)\right\|_{2}^{2}-\left\|g(s_{t},t)-\nabla_{s_{t}}\log q_{t}(s_{t}|s)\right\|_{2}^{2}\right).$
Then, minimizing ^$\hat{\ell}(\pi)$ is equivalent to solving the two-player zero sum game $\min_{\pi}\max_{g}\mathcal{L}(\pi,g).$ To solve this game, we propose following a no-regret strategy over $g$ against a best response over $\pi$ (i.e. a dual algorithm (Swamy et al., 2021)). Specifically, Algorithm 1 applies Follow-the-Leader (Shalev-Shwartz et al., 2012) to optimize $g$ in Line 3, [^4] Note that $\mathcal{L}(\pi,g)$ is a square-loss functional with respect to $g$. Since square loss is strongly convex, FTL is no-regret. In contrast, when optimizing an IPM, one has to use Follow-the-Regularized-Leader (FTRL) to have the no-regret property (e.g., Sun et al. (2019); Swamy et al. (2021)), making implementation more complicated. and performs best response computation over $\pi$ via RL (Line 4) under cost function $c^{k}(s)$: